OpenStudy (anonymous):

solve this equation -9h-6+12h+40=22

OpenStudy (anonymous):

h - 4

OpenStudy (anonymous):

-9h - 6 + 12h + 40 = 22 combine like terms 3h - 6 + 40 = 22 3h - 6 = -18 3h = - 12 divide h = -4

OpenStudy (anonymous):

thank u soooo much